OpenStudy (anonymous):

Find the least common multiple of x³ – x² + x – 1 and x² – 1 . Write the answer in factored form.

OpenStudy (anonymous):

\[x ^{2} – 1=(x-1)(x+1)\] \[x^3 - x^2 + x - 1=(x-1)(x^2+1)\]

OpenStudy (triciaal):

pretend you are doing regular long division and divide first expression by x-1

OpenStudy (anonymous):

@mustafa2014 I'm confused about what least common multiple means, can you explain that please?

OpenStudy (anonymous):

http://www.mathsisfun.com/least-common-multiple.html

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

welcome :)